OpenStudy (anonymous):

What does x equal if 5/x=9/2x

OpenStudy (anonymous):

no value of x satisfies that. because if you rearrange it you would get x=10x/5 as you know no number can be defined like that unless it is 0

OpenStudy (aravindg):

the x terms on both sides gets cancelled leaving away no solution

OpenStudy (anonymous):

That is the solution-- but I am curious why can't you make it x=9/10x? and then go from there 10x^2=9 x^2=9/10 x=3/sqrt10

OpenStudy (aravindg):

5/x=9/2x multiply both sides by x 5=9/2???? no solution

OpenStudy (anonymous):

Why can't you multiply both sides by 1/5 instead of x?

OpenStudy (shubhamsrg):

an unreal solution would be x=infinity infact 2 unreal solutions -> inf and -inf

OpenStudy (hba):

There would be infinite solutions ? @shubhamsrg or no solutions ?

OpenStudy (hba):

I guess they are infinite solutions.

OpenStudy (saifoo.khan):

No sol.

OpenStudy (anonymous):

I understand the answer is no solution--- what I am asking is why can't you divide by 5 to get x by itself.

OpenStudy (shubhamsrg):

2 solns, 2 unreal solutions, or should i say 2 theoretical solutions.

OpenStudy (hba):

I guess we should call it infinite solutions.

OpenStudy (saifoo.khan):

no @hba

OpenStudy (shubhamsrg):

hmm ? I'd still say 2 solutions which are inf and -inf , but both are theoretical .

OpenStudy (hba):

@saifoo.khan Explain. I remember hero telling me that there would be infinite solutions. I also believed there would be no solution.

OpenStudy (anonymous):

Will one of you brilliant people just explain what algebraic rule I am missing/perverting that is making me not see the no solution. Why can't you multiply by 1/5?

OpenStudy (hba):

@sleipnir234 Ok Lets follow you, \[\huge\frac{ 5 }{ x }=\frac{ 9 }{ 2x }\] Now multiply and divide by 1/5. Let us see what you get.

OpenStudy (hba):

Now multiply both sides by 1/5 *

OpenStudy (hba):

\[\frac{ 5 }{ x } \times \frac{ 1 }{ 5 }=\frac{ 9 }{ 2x } \times \frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

x=9/10x

hartnn (hartnn):

if you multiply both sides by 1/5, you get 1/x =9/(10x) and not x=9/10x

hartnn (hartnn):

if there is an identity after cancelling 'x', then there are infinite solutions @hba, here we don't get an identity, ,we get 1=9/10 which is NEVER tru , so, no solution. if we got something like 1=9/9, then we would have infinite solutions

OpenStudy (hba):

@hartnn Thanks for explaining :)

OpenStudy (anonymous):

ohhhhhhhhh Got it. Thank you so much. I'm sorry. It is late. And I am an idiot. Thank you all so much for being patient!!

OpenStudy (shubhamsrg):

There are, 2 solutions! 2 theoretical solutions, one should not ignore that.

OpenStudy (hba):

Welcome :)

hartnn (hartnn):

silly mistakes usually happen,no problem, welcome ^_^